We have one working clock circuit for a cymac main chassis. We have two more PCBs which are both slightly incomplete.
Today I went to Downs and Todd flashed our 3 arduino nanos with the software outlined in LIGO-E2300449. The steps look simple enough, but since Todd had already done this for his Arduinos it was as simple as connecting to his computer and hitting upload.
I then returned to bridge where I tested one of our clock circuits. I learned that the ATX power supply only outputs a voltage if the large ATX connector is plugged in, hence the PCIe backplane resting on the anti-static bag in the photo.
The one complete clock PCB that we have works. We have two more PCBs from Todd, but they are each missing a component or two. One is missing two buffers (center and marked by orange) and the other is missing an RF relay (look top, K1) Perhaps I will work with Todd to complete them.
It is also possible that we don't need the relay, which appears from the schematic to be related to the external input.
The clock circuit which is missing an RF switch (component K1) produces three 65kHz outputs above 5 volts. The external clock input has not been tested.
In order to facilitate the development of a lab power supply standard, I have performed a census of all the devices currently consuming DC power. I have divided the space into three main locations, presumably each location will be served by its own supply.
model | voltage | current | number | |
RF amplifier (AOM driver) | ZHL-1-2W+ | 24V | 0.9A | 4 |
opamp (piezo driver) | EVAL-ADHV4702-1 | 15V | ? | 2 |
model | voltage | current | number | |
Photodiode | PDA**** | +12V, -12V | ? | 9 |
Photodiode | 1811-FS | +15V, -15V | 200mA | 1 |
flipper mount | MFF101 | 15V | ? | 2 |
model | voltage | current | |
Teraxion | Ultra-Narrow DFB Laser Module – LXM | 5V | 1A |
Thorlabs | ULN15TK | 12 V | ? |
Here is a pinout for the AOM Power connector for properly creating the cables using a 5-pin connector.
The pinout is intended for the plug/male type connector, and the cable colors are to match the diagram. The 5-pin connection has two GND and +24V receptacles.
Always double-check the output on each pin before plugging it into an AOM chassis.
NOTE: The D-shaped symbol at the bottom of the pin diagram refers to the connector slot and should be used to orient the connector and how you wire it!
I set up a remote USB hub on the shelf over the optics table. Currently the TeraXion laser is connected, and I successfully sent commands via the control GUI on Brewster. The space under the table is getting a bit cluttered, perhaps we should think of a way to keep these cables off the floor.
I used the 'Mr grip' stripped screw hole thread repair kit and some longer screws to replace a door stop on the door connecting B111B and B111D. I used two strips per hole.
[Jeff, Torrey]
We had previously reported on some high frequency modulations in the error signal while locking the cavities with the piezo and EVAL-ADHV4702-1CPZ amplifier.
This is an opamp with the standard differential input, and I (Jeff) previously created this this splitter doohicky to connect the input ground to the negative port of the amplifier and the input signal to the positive port.
We found that upon removing this from the setup we successfully removed the spurrious pulses from the error signal. Currently the negative input of the amplifier is SMA terminated to the ground of the amplifier, and the input is connected to the positive input.
[Jeff, Sander, Alex, Daniel]
I cleaned the inside of the 5" long, 10" to 8" nipple. We then slid the input vacuum cube on Laser Filter Cavity (LFC) toward the 5 in Long 8 in CF flange bellows. We stuck screws through the nipple and eventually through the bellows as well, getting them somewhat tight with plate nuts. We then further slid the input vacuum cube into place and bolted it down. Jeff and I then removed the screws on the top half of the connection and loosened the others so that we can drop in the copper gasket. The gasket was not properly seated intially, so Jeff pushed the gasket up and away and the gasket dropped into place. We reattached the screws and tightened all the screws. I still need to tighten the screws to ~23 Nm. Waiting a bit for the bellows to plastically deform back might make this easier.
I thought we would want to use the engine hoist, but we decided to just push the cube. We did make some metal to metal contact, but I think the knife edge is ok. I briefly inspected the window attached to the output cube since it is fragile, and it looked fine. I have no real reason to expect damage. I'll install the input window last.
I measured the length from the center to center of the vacuum cubes (around where the mirrors will go), and it is 44 in (1.12 in), almost exactly what Ian designed.
See attached photo.
I tightened the connection to 20 Nm. There is a stainless to stainless connection around ~90% of the flange surface.
I loaned ~15 1.1 Ω resistors (they need a 0.1 Ω resistor) to the Hutzler Lab from the rotating electronics rack from the Kimble Lab in B102B. Point of contact is Yuiki Takahashi.
[Jeff, Daniel]
I cleaned one of the 8" CF Flanges on a 5" long bellows from the Holometer from tape and tape residue. I then cleaned the other end and Jeff and I installed it on the open end of the reducing tee we just installed. The bolts have to be installed from the tee because of the ripple pattern on the bellows. The 2" long bolts can be inserted along side the 6" flange on the tee. I tightened the bolts to 23 Nm to get a steel to steel contact, 2 Nm more than the last 8" CF Thru flange connection.
I thought we could use a 6 way reducer cross with a 6" CF Flange (qty. 2) to 2.75 in CF Flange (qty. 4) from Alan Rice, but one of the knife edges is damaged. It was on a blank CF Flange, so maybe vacuum was still able to be pulled. I would rather not have to figure this out after cleaning and installing the part, so I will not use it for now. See attached photos of the damage.
[Jeff, Daniel]
We removed the tape and cleaned the residue from one of the 8" CF Flanges on a 8 in to 6 in Reducer Tee from the Holometer. We then attached it to the 10 in CF to 8 in CF 12 in Long Custom Nipple on the Laser Filter Cavity (LFC) output vacuum cube. We tried to make the 6" CF Flange stick out vertically as best as possible since the 8" flange on the custom nipple is rotatable. The 8" flange on the nipple on the input vacuum cube is also rotatable, so we should be able to align the bolt holes without being constrained by this. I started to implement the 80-20 structure to support the reducer tee. The bolts are quite bad the hex input is easily stripped. We should get better screws (they are 1/4-20 button head socket cap screws, around 1/2" length but I need to check). We have not yet tightened the bolts on the CF Connection.
The 80/20 bolts are speced to be and are 1/2" long.
I tightened the bolts on the CF connection to 20 Nm. The flanges were pretty flush all the way around with no copper visible.
Next, we need to clean one of the bellows' flanges (the other should be clean) and attach it to the tee. We then need to slide the cube into the bellows. Maybe we could use the engine hoist to remove some of the weight of the cube.
[Sander, Alex, Daniel]
We acquired some vacuum components from Alan Rice from the subbasement of Kellogg. All of the vacuum equipment is dirty so would need to be cleaned and baked, but much of it could be used in our experiments. See this sheet for an inventory. Note that some components were in a beam line and could have been radioative. Alan measured every component with a geiger counter to ensure that they are not radioactive.
When zooming in to perform a ringdown measurement I found a ~200 MHz oscillation in the TRANS signal. This is OFC2 TRANS. The oscillation frequency corresponds exactly to the opposite AOM's frequency. Previously we had observed a beatnote between the two AOM frequencies, and could be solved by tuning the AOM frequencies to be more than 2 ish MHz away from each other.
Now we are seeing a 200 MHz signal regardless of the relative frequency. Changing the other AOM frequency changes this oscillation directly. I don't think this is optical, and is purely an electrical pick up. Just making a note of this to be seen on the log book for the future. Also this is a one way problem, OFC2 sees AOM1 but OFC1 does not see AOM2.
I added 1.75" long 5/16-24 Screws, Washers, and Nuts to the Agilent TwisTorr 74 Turbo Vacuum Pump to a CF 4.5" Flange to KF 50 adaptor. Neither flange nuts nor washers fit next to the turbo pump, so I just used regular nuts. We still need to tighten the screws. 1.75" long screws were the perfect length for this application (matching the length of the nuts), but 2" screws, which are standard for connecting 4.5" CF Flanges with through holes, would also work.
I tightened the bolts to 34 Nm (incrementing by 3 Nm from 13 Nm) with a 1/2 open faced wrench holding the nuts. Tightening the bolts without the wrench did not actually compress the copper gasket. The gasket was still a bit visable, but it will be easy to tighten further if needed.
It looks like the manual recommends 20 Nm: https://www.agilent.com/cs/library/usermanuals/public/TwisTorr%2074%20FS.pdf
I removed the duct tape and duct tape residue from the outside of the 8 in CF Flange portion of the 10 in CF Flange to Rotatable 8 in CF Flange 5 in Long on the Laser Filter Cavity (LFC) Input Vacuum Cube. I used clean room wipes with acetone and isopropanol. I wrapped the open end in UHV Foil and moved the cube into place. We should do a quick clean of the inside of the nipple before final installation.
I then placed the 80-20 structure on the optics table. I will add the last pieces tomorrow.
[Alex, Daniel]
I got two aluminum KF50 Centering Rings from Nick Hutzler's group and machined them in the lathe to remove one of their lips so that they can hold a PCB that interfaces with the inside and outside of the Dewer. I clamped them on the outside lip with a 6 jaw chuck. I used a moderate amount of clamping force (~90 degrees of rotation with the chuck key) to hold them. I slowly increased the spindle speed to ensure the rings wouldn't fly off; I used a final spindle speed of 1700 rpm, which is pretty standard for aluminum and a carbide tool. I kept the x-axis of the lathe at around 1.93" and slowly moved the tool in the z-axis to remove the lip. I needed to take off 0.080" of material. For the last few thousands of an inch, I moved the tool in the x-axis instead of the z-axis. I then used a deburring tool by hand to remove any burrs from machining. I could not see any markings from the clamping jaws.
Tomorrow, I will clean and start to bake out these centering rings.
We cleaned the centering rings with the normal aluminum procedure of successive 3 minute baths and scrubs of 1:30 Simple Green:DI water, DI water, and isopropanol. We put it in the vacuum oven to bake out at 120° for 48 hours since we don't know the alloy of aluminum.
I turned off the vacuum pump and closed the valve to the vacuum pump around noon today. According to the vacuum oven, the bakeout lasted 49-50 hours. We can store these parts under vacuum unitl we need them or the vacuum oven.
I machined 3 more aluminum KF50 centering rings exactly like I did the first 2.
We moved the two cleaned centering rings out of the oven and put them in UHV foil and an antistatic bag. We repeated the cleaning process for the three dirty rings and are baking them out.
I turned off the vacuum pump and closed the valve to the vacuum pump